Experiment status

From Hall A Wiki
Jump to: navigation, search

Experiment down while ESR is recovering ( should be back up around midnight )

  • Magnets for spectrometers need to be cooled back down ( after Hall B minitorus cooldown ) 6 hours
  • Target He3 needs to be cooled down 6 hours
  • AC of the left arm needs to be hooked up ( can be done in parallel with cooldown )

-- 19:42, May 3, 2007 (EDT)